# spacing=nonfrench

### v3

@@ 8. Conclusion
This paper introduces the first smart surface-enabled 5G mmWave roadside networking system designed to efficiently steer outdoor mmWave signals into vehicles, enable simultaneous data exchange and neighboring cell measurements and reliable make-before-break handovers. We conduct an extensive evaluation in various indoor and outdoor multi-user settings, demonstrating significant TCP improvement. Future work will explore Wall-Street’s performance under various environmental conditions and mobility scenarios, including different weather and terrain variations.
@@ Appendix A Surface Design Parameters
In this section, we provide design parameters of Wall-Street’s electric and magnetic meta-atom, as illustrated in Fig.˜22. The thickness of PCB Roger 4003C substrate is $0.406$ mm and the thickness of the copper trace is $1$ oz. To simulate the unit cell response, we create an air box that defines the boundary conditions. The top and bottom boundaries of the box are perfect electric boundary, and the front and back of the box are perfect magnetic boundary.
Appendix B Beam Angle Selection
We conduct Wall-Street beam pattern microbenchmarks using angle pairs of -40°/40°, -40°/55°, and -68°/45°. These angles reflect typical mmWave handover scenarios, where handovers occur every 0.13 km within a gNB’s typical coverage radius of 0.15 km (Hassan et al., 2022). The demonstration video of different multi-beam angles are also available in bit.ly/42xb2nr.
Figure 23. Beam angle selection for microbenchmark.

### v4

@@ 8. Conclusion
This paper introduces Wall-Street that efficiently steers outdoor mmWave signals into vehicles, simultaneously exchange data and and scan neighboring cells, and realizes make-before-break HOs. We integrated the surface into the COSMOS testbed and conducted driving experiments with surface-mounted SUV and multiple gNB/UE nodes. Our trace-driven ns-3 simulations show the TCP improvement by up to 78% and latency reduction by up to 34%.
Acknowledgement
This material is based upon work supported by the National Science Foundation under Grant Nos. CNS-2148271, CNS-2433915, CNS-2450567 and is supported in part by funds from federal agency and industry partners as specified in the Resilient & Intelligent NextG Systems (RINGS) program. We gratefully acknowledge the support of a grant from the Princeton University School of Engineering and Applied Science Innovation Fund.
(a) Electric meta-atom.
(b) Magnetic meta-atom.
Figure 21. Wall-Street’s unit cell design parameters.
@@ Appendix A Surface Design Parameters
In this section, we provide design parameters of Wall-Street’s electric and magnetic meta-atom, as illustrated in Fig.˜21.
Appendix B Link Budget Calculation
To demonstrate the deployment feasibility, we perform a link-budget analysis using a realistic in-vehicle deployment scenario. First, we calculate the downlink SNR at the in-vehicle UE under worst-case conditions:
(9) $SNR_{UE}~(dB)=P_{gNB}-L_{gNB}-L_{window}+G_{RIS,Rx}+$
$G_{RIS,Tx}-L_{UE}+G_{UE}-P_{nf}$
[…]
We also evaluate the feasibility of gNB-to-gNB measurement, where the serving gNB measures a signal from the neighbor gNB reflected by the in-vehicle RIS. We calculate the reflective SNR measured at the serving gNB under worst-case conditions as:
(10) $SNR_{gNB}~(dB)=P_{gNB_{n}}-L_{gNB_{n}}-L_{window}+G_{RIS,Rx}+$
$G_{RIS,Tx}-L_{window}-L_{gNB_{s}}+G_{gNB_{s}}-P_{nf}$
[…]